17. Given the following supply and demand equations

QD = 100 – 5P

QS = 10 + 5P

a) Determine the equilibrium price and quantity.

b) If the government sets a minimum price of $10 per unit, how many units would be supplied and how many would be demanded?

c) If the govt. sets a maximum price of $5 per unit, how many units would be supplied and how many would be demanded?

d) If the demand increases to

Q’D = 200 – 5P

Determine the new equilibrium price and quantity.
1
Expert's answer
2020-12-01T10:26:57-0500

a) The equilibrium is at

"Q_d = Q_s \\

100 – 5P = 10 + 5P \\

P=9 \\

Q = 100 - 5\times 9 = 55"

b)

"P=10 \\

Q_d=100-5 \times10 = 50 \\

Q_s=10+5 \times 10= 60"

The surplus = 60 - 50 = 10 units

It will maintain it by using law and order

c)

"P=5 \\

Q_d= 100 - 5\times 5 = 75 \\

Q_s=10+5 \times 5 = 35 \\"

The shortage =75 – 35 = 40 units

d) New equilibrium

"200 - 5P = 10+5P \\

P=19 \\

Q=10+5 \times 19 \\

Q =105"